OpenStudy (ledah):

The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (anonymous):

well volume = a b c new volume = 3a 3b 3c

OpenStudy (anonymous):

3*3*2=27 so newarea=27*a b c

OpenStudy (anonymous):

oops i mean newvolume new volume will be 27 times bigger :O

OpenStudy (ledah):

Don't really get what you're saying

OpenStudy (anonymous):

volume=a*b*c a b and c are the sides if you triple each side then new volume=(3*a) * (3*b) * (3*c) this can be rewritten using algebra as newvolume=27 * (a*b*c) ( ) notice abc is your original volume. so if you triple each side then the new volume = original volume *27

OpenStudy (ledah):

So I have to find the original volume?

OpenStudy (ledah):

is that what I have to do? or do I just multiply 150*3